OpenStudy (anonymous):

f(x)=x^3-x^2-36x+36/x^3-17x^2+90x-144, f(x) has one hole point at(__,__)

OpenStudy (amistre64):

someone really wants you to learn to factor

OpenStudy (anonymous):

well i know how to factor that out but i dont know what the hole point is

OpenStudy (amistre64):

the hole appears when the top and bottom have a like factor

OpenStudy (amistre64):

determine the common roots and youll see your "x" value; simplify and input the x again to get the y

OpenStudy (anonymous):

never mind this one doesnt seem like it factors nicely

OpenStudy (anonymous):

it will factor, find 1 rational factor and then use synthetic division to get the rest

OpenStudy (amistre64):

for example: \[\frac{(x-1)(x+2)}{(x+2)(x-3)}\] x+2 is common and has a root at -2 \[\frac{(-2-1)\cancel{(x+2)}}{\cancel{(x+2)}(-2-3)}=\frac{3}{5} \]

OpenStudy (anonymous):

even easier way is to use grouping method on first fraction

OpenStudy (anonymous):

numeratpr(

OpenStudy (anonymous):

i tried factoring but cant do i have to use the quadratic formula or am i just making it harder than what it is

OpenStudy (anonymous):

x^3-x^2-36x+36 x^2(x-1)-36(x-1) (x-1)(x^2-36) (x-1)(x-6)(x+6) <===numerator

OpenStudy (accessdenied):

Quadratic formula is for quadratics, not cubics

OpenStudy (anonymous):

lol "someone really wants you to learn to factor" And this is why high school math before AP sucked.

OpenStudy (anonymous):

For the denominator, i'm going to find a zero on a calculator and use that to try synthetic division

OpenStudy (anonymous):

ok thanks cause i get for the denominator x^2(x-17)18(5x-8) and that cant be right

OpenStudy (anonymous):

x^3-17x^2+90x-144, 3| 1 -17 90 -144 3 -42 144 --------------------------- 1 -14 48 | 0 x^2-14x+48=0 (x-6)(x-8)=0 denominator is (x-3)(x-6)(x-8) So in factored form numerator/denominator: \[\frac{(x-1)(x-6)(x+6)}{(x-3)(x-6)(x-8)}=\frac{(x-1)(x+6)}{(x-3)(x-8)}\]

OpenStudy (anonymous):

ok thanks for doing that but how do i find the hole

OpenStudy (accessdenied):

to quote amistre64: "the hole appears when the top and bottom have a like factor" There is one set of like factors in the numerator and denominator: x - 6 So, there would be a hole at x - 6 = 0, or x = 6

OpenStudy (anonymous):

i've actually got 2 holes

OpenStudy (anonymous):

one at x=3, x=8

OpenStudy (anonymous):

actually, x=3,x=8 are vertical asymptotes, not holes

OpenStudy (anonymous):

x=6 would be the hole

OpenStudy (anonymous):

so its 6,0 and you got the six becasue thats what was common in both of them and you canceled it

OpenStudy (anonymous):

f(x) is undefined at x=3,6,8. But because (x-6) is also a factor in the numerator, it is only a hole in the graph, not a vertical asymptote

OpenStudy (anonymous):

ok im just trying to figure out how to write it cause it gives me 2 spaces to enter a number and the first space is 6 but i cant fiugure out what goes in the next space

OpenStudy (anonymous):

They want a y-value then?

OpenStudy (anonymous):

\[\frac{(x-1)(x+6)}{(x-3)(x-8)}\], plug 6 into this expression i posted above

OpenStudy (anonymous):

You get the point (6,-10)

OpenStudy (anonymous):

ok got it thanks so much